OpenStudy (anonymous):

√6 multiplied by (7√3 +6)

OpenStudy (unklerhaukus):

\[√6 \times (7√3+6)\]\[=√6\times 7√3+√6 \times 6\]\[=7√18+6√6\]

OpenStudy (anonymous):

Can it be simplified further?

OpenStudy (unklerhaukus):

\[=21√2+6√6\]

OpenStudy (anonymous):

Multiply it out you get: 7√18+6√6 = 7√3*3*2 + 6√6 = 21√2 + 6√6

OpenStudy (anonymous):

Thank you!

OpenStudy (unklerhaukus):

\[=(21+6√3)√2\]